Show ALL work to receive rating. Thanks!

1. Let f(x) = -4x^3+6x^2

a) Where is f(x) increasing/decreasing? Make a sign chart.

b) Classify the critical points as local max, local min, or neither.

c) Where is f(x) concave up/concave down? Does it have any points of inflection?

d) Use the information above to sketch the curve. Note that f(1/2) = 1. Be sure your graph includes the x and y intercepts if they exist.
Answer #1

giren that f(x) = - 4x3 + 6x² so fix) = -127?+ 12x for critical points put f(X=0 - 12x ² + 124 = 0 2-2:0 ² x(x-1) = 0 x=0 o